OpenStudy (anonymous):

The sum of 3 consecutive odd integers is -21. Find the "middle" number. ?

OpenStudy (lastdaywork):

Also post your attempt. This will allow OS community to know exactly where you need help.

OpenStudy (anonymous):

It's -7.

OpenStudy (anonymous):

Because -6 + -7 + -8 = -21!

OpenStudy (anonymous):

thanks @xxxkittyxxx (:

OpenStudy (anonymous):

You're welcome!

OpenStudy (texaschic101):

x = 1st integer x + 2 = 2nd integer x + 4 = 3rd integer x + x + 2 + x + 4 = -21 -- combine like terms 3x + 6 = -21 3x = -21 - 6 3x = - 27 x = -9 x + 2 = -9 + 2 = -7 x + 4 = -9 + 4 = -5 your numbers are : -5,-7,-9 your middle number is -7